OpenStudy (anonymous):

*** Will give you a medal *** Compare the graphs of the inverse variations by comparing asymptotes, similar points, lines of reflections, and quadrants. y=-0.7/x and y=-0.9/x

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

are you able to graph these two equations

OpenStudy (anonymous):

yeahh

jimthompson5910 (jim_thompson5910):

what similarities do you see

OpenStudy (anonymous):

they are both negatives and decimals

jimthompson5910 (jim_thompson5910):

good, so they are reflected versions (over the x-axis) of y = 1/x

OpenStudy (anonymous):

yes.

jimthompson5910 (jim_thompson5910):

they are also compressed a bit compared to y = 1/x, although, y = -0.7/x is a bit more compressed than y = -0.9/x

jimthompson5910 (jim_thompson5910):

which quadrants do both go through?

OpenStudy (anonymous):

3 and 4 ?

jimthompson5910 (jim_thompson5910):

close, but it's 2 and 4

jimthompson5910 (jim_thompson5910):

and finally, what are the asymptotes of each

OpenStudy (anonymous):

i have no idea.. idk how to do those.

jimthompson5910 (jim_thompson5910):

what's the vertical asymptote of y = 1/x

OpenStudy (anonymous):

idk. idk how to do that.

jimthompson5910 (jim_thompson5910):

notice how if x = 0, then you get an undefined error

jimthompson5910 (jim_thompson5910):

because you can't divide by zero

jimthompson5910 (jim_thompson5910):

so the vertical asymptote is x = 0

OpenStudy (anonymous):

yeah

jimthompson5910 (jim_thompson5910):

the vertical asymptotes of y=-0.7/x and y=-0.9/x are both x = 0 because of that same reason

jimthompson5910 (jim_thompson5910):

similarly, the horizontal asymptotes are y = 0 a graph will help see this and you can find the domain of the inverse to get the same thing

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

thanks (:

jimthompson5910 (jim_thompson5910):

yw